OpenStudy (kingskurt):

Trevor has an investment worth $6,774.50. He made his only deposit in it 22 years ago. Given that the investment yields 2.7% simple interest annually, how big was the initial deposit?

OpenStudy (hayisforhorses):

x*1.027^22=6,774.50 x*1.79700695983=6,774.50 x=6,774.50/1.79700695983 x=3,769.88